# Latent Functional Maps

### v2

@@ Abstract
Neural models learn data representations that lie on low-dimensional manifolds, yet modeling the relation between these representational spaces is an ongoing challenge. By integrating spectral geometry principles into neural modeling, we show that this problem can be better addressed in the functional domain, mitigating complexity, while enhancing interpretability and performances on downstream tasks. To this end, we introduce a multi-purpose framework to the representation learning community, which allows to: (i) compare different spaces in an interpretable way and measure their intrinsic similarity; (ii) find correspondences between them, both in unsupervised and weakly supervised settings, and (iii) to effectively transfer representations between distinct spaces. We validate our framework on various applications, ranging from stitching to retrieval tasks, demonstrating that latent functional maps can serve as a swiss-army knife for representation alignment.
@@ 1 Introduction
Neural Networks (NNs) learn to represent high-dimensional data as elements of lower-dimensional latent spaces. While much attention is given to the model’s output for tasks such as classification, generation, reconstruction, or denoising, understanding the internal representations and their geometry is equally important. This understanding facilitates reusing representations for downstream tasks [48] and the comparison between different models [14, 29], thereby broadening the applicability of NNs, and understanding their structure and properties. Moreover, recent studies have shown that distinct neural models often develop similar representations when exposed to similar stimuli, a phenomenon observed in both biological [9, 16] and artificial settings [28, 14, 29]. Notably, even with different initializations of the same neural architecture, internal representations of distinct models can often be aligned through a linear transformation [47, 38]. This indicates a level of consistency in how NNs process information, highlighting the importance of exploring and understanding these internal representations.
In this paper, we shift our focus from characterizing relationships between samples in distinct latent spaces to modeling a map among function spaces defined on these latent manifolds. To this end, we propose leveraging spectral geometry principles by applying the framework of functional maps [34] to the field of representation learning. Functional maps represent correspondences between function spaces on different manifolds, providing a new perspective on the problem of latent space communication. In this setting, many difficult constraints can be easily manipulated and expressed compactly [35]. For instance, as shown in Figure 1, the mapping in the functional space ($C$) becomes a linear map with a sparse structure.
@@
Figure 1: Framework overview: given two spaces $\mathcal{X}$, $\mathcal{Y}$ their samples lie on two manifolds $\mathcal{M}$, $\mathcal{N}$, which can be approximated with the KNN graphs $G_{\mathcal{X}}$,$G_{\mathcal{Y}}$; We can optimize for a latent functional map $C$ between the eigenbases of operators defined on the graphs. This map serves as a map between functions defined on the two manifolds and can be leveraged for (i) comparing representational spaces, (ii) solving correspondence problems, and (iii) transferring information between the spaces.
@@ 2 Related Work
Similarity between latent spaces
Comparing representations learned by neural models is of fundamental importance for a diversity of tasks, ranging from representation analysis to latent space alignment and neural dynamics. In order to do so, a similarity measure between different spaces must be defined [13]. This can range from functional similarity (matching the performance of two models) to similarity defined in representational space [14], which is where our framework falls in. A classical statistical method is Canonical Correlation Analysis (CCA) [11], known for its invariance to linear transformations. Various adaptations of CCA aim to enhance robustness, such as Singular Vector Canonical Correlation Analysis (SVCCA) [37], or to decrease sensitivity to perturbations using methods like Projection-Weighted Canonical Correlation Analysis (PWCCA) [28]. Closely related to these approaches, Centered Kernel Alignment (CKA) [14] measures the similarity between latent spaces while ignoring orthogonal transformations. However, recent research [7] reveals that CKA is sensitive to local shifts in the latent space.
@@
Latent communication
This relatively new concept, introduced by [29], builds on the hypothesis that latent spaces across neural networks (pre-)trained with many variation factors, from random seed initialization to architecture or even data modality, are intrinsically compatible. This notion is supported by numerous empirical studies [28, 20, 14, 4, 42, 1, 45, 18, 19, 2, 30, 6], with the phenomenon being particularly evident in large and wide models [39, 23]. The core idea is that relations between data points (i.e., distances according to some metric) are preserved across different spaces because the high-level semantics of the data are the same and neural networks learn to encode them similarly [12]. With this "relative representation", the authors show that it is possible to stitch [19] together model components coming from different models, with little to no additional training, as long as a partial correspondence of the spaces involved is known. Indeed, [17, 22, 25, 27] show that a simple linear transformation is usually enough to effectively map one latent space into another, measuring the mapping performance via desired downstream tasks.
@@ 3 Background
Consider two manifolds $\mathcal{M}$ and $\mathcal{N}$ equipped with a basis such that any function $f:\mathcal{M}\rightarrow\mathbb{R}$ can be represented as a linear combination of basis functions $\mathbf{\Phi}_{\mathcal{M}}$: $f=\sum_{i}a_{i}\Phi_{i}^{\mathcal{M}}=\mathbf{a}\mathbf{\Phi}_{\mathcal{M}}$. Given the correspondence $T:\mathcal{M}\rightarrow\mathcal{N}$ between points on these manifolds, for any real-valued function $f:\mathcal{M}\rightarrow\mathbb{R}$, one can construct a corresponding function $g:\mathcal{N}\rightarrow\mathbb{R}$ such that $g=f\circ T^{-1}$. In other words, the correspondence $T$ defines a mapping between two function spaces $T_{F}:\mathcal{F}(\mathcal{M},\mathbb{R})\rightarrow\mathcal{F}(\mathcal{N},\mathbb{R})$. [34] showed how such a mapping is linear and can be represented as a (possibly infinite) matrix $\mathbf{C}$ such that for any function $f$ represented as a vector of coefficients $\mathbf{a}$, we have $T_{F}(\mathbf{a})=\mathbf{C}\mathbf{a}$.
The functional representation is particularly well-suited for map inference (i.e., constrained optimization). When the underlying map $T$ (and by extension the matrix $\mathbf{C}$) is unknown, many natural constraints on the map become linear constraints in its functional representation. In practice, the simplest method for recovering an unknown functional map is to solve the following optimization problem:
@@
where $\mathbf{A}$ and $\mathbf{B}$ are sets of corresponding functions expressed in the bases on $\mathcal{M}$ and $\mathcal{N}$, respectively, and $\rho(\mathbf{C})$ represents additional constraints deriving from the properties of the matrix $\mathbf{C}$ [35]. When the shapes are approximately isometric and the descriptors are well-preserved by the (unknown) map, this procedure provides a good approximation of the underlying map. In the case where the correspondence $T$ is encoded in a matrix $\mathbf{S}$, the functional map can be retrieved as $\mathbf{C}=\mathbf{\Phi}_{\mathcal{N}}^{{\dagger}}\mathbf{S}\mathbf{\Phi}_{\mathcal{M}}$ where $\mathbf{\Phi}_{\mathcal{M}}$ and $\mathbf{\Phi}_{\mathcal{N}}$ are the bases of the functional spaces $\mathcal{F}(\mathcal{M},\mathbb{R})$ and $\mathcal{F}(\mathcal{N},\mathbb{R})$, respectively, and denotes the pseudo-inverse.
4 Method
4.1 Setting
We consider deep neural networks $f:=f_{1}\circ f_{2}\circ...f_{n}$ where each layer $f_{i}$ is associated to a representational space $\mathcal{X}$ corresponding to the image of $f_{i}$. We assume that elements $x\in\mathcal{X}$ are sampled from a latent manifold $\mathcal{M}$. Considering pairs of spaces $(\mathcal{X},\mathcal{Y})$, and corresponding manifolds $\mathcal{M},\mathcal{N}$ our objective is to characterize the relation between them by mapping the space of functions $\mathcal{F}(\mathcal{M})$ to $\mathcal{F}(\mathcal{N})$. Our framework is depicted in Figure 1. In the following, we will start by approximating $\mathcal{X}$ from a sample estimate, building a graph in the latent space.
Building the graph
To leverage the geometry of the underlying manifold, we model the latent space of a neural network building a symmetric k-nearest neighbor (k-NN) graph [21]. Given a set of samples $X=\{x_{1},\ldots,x_{n}\}$, we construct an undirected weighted graph $G=(X,E,\mathbf{W})$ with nodes $X$, edges $E$, and weight matrix $\mathbf{W}$. The weight matrix is totally characterized by the choice of distance function $d(x_{i},x_{j})$ with $x_{i},x_{j}\in X$. Suitable choices include the L2 metric or the angular distance. Edges $E$ are defined as $E=\{(x_{i},x_{j})\in X\times X\mid x_{i}\sim_{\text{k}}x_{j}\text{ or }x_{j}\sim_{\text{k}}x_{i}\}$, where $x_{i}\sim_{\text{k}}x_{j}$ indicates that $x_{j}$ is among the $k$ nearest neighbors of $x_{i}$. The weight matrix $\mathbf{W}\in\mathbb{R}^{n\times n}_{\geq 0}$ assigns a weight $\omega(x_{i},x_{j})$ to each edge $(x_{i},x_{j})\in E$, and $\mathbf{W}_{i,j}=0$ otherwise.
Next, we define the associated weighted graph Laplacian $\mathcal{L}_{G}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $\mathbf{D}_{i,i}=\sum_{j=1}^{n}\mathbf{W}_{i,j}$. $\mathcal{L}_{G}$ is a positive semi-definite, self-adjoint operator [44]), therefore, it admits an eigendecomposition $\mathcal{L}_{G}=\mathbf{\Phi}_{G}\mathbf{\Lambda}_{G}\mathbf{\Phi}_{G}^{T}$, where $\mathbf{\Lambda}_{G}$ is a diagonal matrix containing the eigenvalues, and $\mathbf{\Phi}_{G}$ is a matrix whose columns are the corresponding eigenvectors. The eigenvectors form an orthonormal basis for the space of functions defined on the graph nodes (i.e., $\mathbf{\Phi}_{G}^{T}\mathbf{\Phi}_{G}=\mathbf{I}$).
Throughout this paper, we assume the eigenvalues (and corresponding eigenvectors) are sorted in non-descending order $0=\Lambda_{1}\leq\Lambda_{2}\leq\dots\leq\Lambda_{n}$. One may consider a subset of eigenvectors, namely those associated with the $k$ smallest eigenvalues, to compactly approximate a graph signal, employing techniques akin to Fourier analysis.
As demonstrated in many recent works [41, 5], the eigenvalues and eigenvectors of the graph Laplacian associated with a k-NN graph approximate the weighted Laplace-Beltrami operator, placing us in a setting similar to the original one of [34].
@@ 4.2 Latent Functional Maps
We consider pairs of latent spaces $\mathcal{X}$ and $\mathcal{Y}$. We model each space using a subset of training samples $X=\{x_{1},\ldots,x_{n}\}$ and $Y=\{y_{1},\ldots,y_{n}\}$ and build the k-NN graphs $G_{X}$ and $G_{Y}$ from these samples, respectively. For each graph, we compute the graph Laplacian $\mathcal{L}_{G}$ and derive the first $k$ eigenvalues $\Lambda_{G}$ and eigenvectors $\mathbf{\Phi}_{G}=[\phi_{1},\ldots,\phi_{k}]$, which serve as the basis for the function space defined on the latent spaces.
@@
Given the set of corresponding functions $\mathbf{F}_{G_{X}}=[f^{G_{X}}_{1},\ldots,f^{G_{X}}_{n_{f}}]$ and $\mathbf{F}_{G_{Y}}=[f^{G_{Y}}_{1},\ldots,f^{G_{Y}}_{n_{f}}]$, we consider the optimization problem defined in Equation 1 and incorporate regularizers for Laplacian and descriptor operator commutativity, as defined in [31]:
$\underset{\mathbf{C}}{\arg\min}||\mathbf{C}\hat{\mathbf{F}}_{G_{X}}-\hat{\mathbf{F}}_{G_{Y}}||^{2}+\alpha\rho_{\mathcal{L}}(\mathbf{C})+\beta\rho_{f}(\mathbf{C})$ (2)
where $\hat{\mathbf{F}}_{G}=\mathbf{\Phi}_{G}^{T}\mathbf{F}_{G}$ are the spectral coefficients of the functions $\mathbf{F}_{G}$, $\rho_{\mathcal{L}}$ and $\rho_{f}$ are the Laplacian and descriptor operator commutativity regularizers respectively. We specify how we compute the regularizers in Appendix A.
As a set of corresponding functions, we use the geodesic distance functions computed from a point $x\in X$ to all other points in $X$, where $x$ is a point for which we know the corresponding point $y\in Y$ in the other latent space. Once we have solved the optimization problem defined in Equation 2, we refine the resulting functional map $\mathbf{C}$ using the algorithm proposed by [24].

### v3

@@ Abstract
Neural models learn data representations that lie on low-dimensional manifolds, yet modeling the relation between these representational spaces is an ongoing challenge. By integrating spectral geometry principles into neural modeling, we show that this problem can be better addressed in the functional domain, mitigating complexity, while enhancing interpretability and performances on downstream tasks. To this end, we introduce a multi-purpose framework to the representation learning community, which allows to: (i) compare different spaces in an interpretable way and measure their intrinsic similarity; (ii) find correspondences between them, both in unsupervised and weakly supervised settings, and (iii) to effectively transfer representations between distinct spaces. We validate our framework on various applications, ranging from stitching to retrieval tasks, and on multiple modalities, demonstrating that Latent Functional Maps can serve as a swiss-army knife for representation alignment.
@@ 1 Introduction
Neural Networks (NNs) learn to represent high-dimensional data as elements of lower-dimensional latent spaces. While much attention is given to the model’s output for tasks such as classification, generation, reconstruction, or denoising, understanding the internal representations and their geometry is equally important. This understanding facilitates reusing representations for downstream tasks [61, 14] and the comparison between different models [23, 39], thereby broadening the applicability of NNs, and understanding their structure and properties. Moreover, recent studies have shown that distinct neural models often develop similar representations when exposed to similar stimuli, a phenomenon observed in both biological [16, 25, 62] and artificial settings [38, 23, 39]. Notably, even when neural networks have different architectures, internal representations of distinct models can often be aligned through a linear transformation [60, 49]. This indicates a level of consistency in how NNs process information, highlighting the importance of exploring and understanding these internal representations.
In this paper, we shift our focus from characterizing relationships between samples in distinct latent spaces to modeling a map among function spaces defined on these representational spaces. To this end, we propose leveraging spectral geometry principles by applying the framework of functional maps [44] to the field of representation learning. Functional maps represent correspondences between function spaces on different manifolds, providing a new perspective on the problem of representational alignment. In this setting, many complex constraints can be easily expressed compactly [45]. For instance, as shown in Figure 1, the mapping ($C$) between two spaces $\mathcal{M}$ and $\mathcal{N}$ can be represented in the functional space as a linear map with a sparse structure.
@@
Figure 1: Framework overview: given two spaces $\mathcal{X}$, $\mathcal{Y}$ their samples lie on two manifolds $\mathcal{M}$, $\mathcal{N}$, which can be approximated with the KNN graphs $G_{\mathcal{X}}$,$G_{\mathcal{Y}}$. We can optimize for a latent functional map $C$ between the eigenbases of operators defined on the graphs. This map serves as a map between functions defined on the two manifolds and can be leveraged for (i) comparing representational spaces, (ii) solving correspondence problems, and (iii) transferring information between the spaces.
@@ 2 Related Work
Similarity between latent spaces.
Comparing representations learned by neural models is of fundamental importance for a diversity of tasks [62], ranging from representation analysis to latent space alignment and neural dynamics. In order to do so, a similarity measure between different spaces must be defined [22]. This can range from functional similarity (matching the performance of two models) to similarity defined in representational space [23], which is where our framework falls in. A classical statistical method is Canonical Correlation Analysis (CCA) [18], known for its invariance to linear transformations. Various adaptations of CCA aim to enhance robustness, such as Singular Vector Canonical Correlation Analysis (SVCCA) [47], or to decrease sensitivity to perturbations using methods like Projection-Weighted Canonical Correlation Analysis (PWCCA) [38]. Closely related to these approaches, Centered Kernel Alignment (CKA) [23] measures the similarity between latent spaces while ignoring orthogonal transformations. However, recent research [10] reveals that CKA is sensitive to local shifts in the latent space.
@@
Representation alignment.
Complementary to measuring the similarity of distinct representational spaces, several methods have been proposed to optimize for a transformation to align them [1]. The concept of latent space communication, introduced by [39], builds on the hypothesis that latent spaces across neural networks, varying random seed initialization to architecture or even data modality, are intrinsically compatible. This notion is supported by numerous empirical studies [38, 30, 23, 6, 54, 3, 57, 27, 29, 4, 40, 9, 15, 8], with the phenomenon being particularly evident in large and wide models [50, 33]. The core idea is that relations between data points (i.e., distances according to some metric) are preserved across different spaces because the high-level semantics of the data are the same and neural networks learn to encode them similarly [19]. With this "relative representation", the authors show that it is possible to stitch [29] together model components coming from different models, with little to no additional training, as long as a partial correspondence of the spaces involved is known. Indeed, [26, 32, 35, 37] show that a simple linear transformation is usually enough to effectively map one latent space into another, measuring the mapping performance via desired downstream tasks.
@@ 3.1 Background
Consider two manifolds $\mathcal{M}$ and $\mathcal{N}$ equipped with a basis $\mathbf{\Phi}^{\mathcal{M}}$ (respectively $\mathbf{\Phi}^{\mathcal{N}}$). Any squared integrable function $f:\mathcal{M}\rightarrow\mathbb{R}$ can be represented as a linear combination of basis functions $\mathbf{\Phi}_{\mathcal{M}}$: $f=\sum_{i}a_{i}\Phi_{i}^{\mathcal{M}}=\mathbf{a}\mathbf{\Phi}_{\mathcal{M}}$. Given a bijective correspondence $T:\mathcal{M}\rightarrow\mathcal{N}$ between points on these manifolds, for any real-valued function $f:\mathcal{M}\rightarrow\mathbb{R}$, one can construct a corresponding function $g:\mathcal{N}\rightarrow\mathbb{R}$ such that $g=f\circ T^{-1}$. In other words, the correspondence $T$ defines a mapping between two function spaces $T_{F}:\mathcal{F}(\mathcal{M},\mathbb{R})\rightarrow\mathcal{F}(\mathcal{N},\mathbb{R})$. Such a mapping is linear [44] and can be represented as a matrix $\mathbf{C}$ such that for any function $f$ represented as a vector of coefficients $\mathbf{a}$, we have $T_{F}(\mathbf{a})=\mathbf{C}\mathbf{a}$.
The functional representation is particularly well-suited for map inference (i.e., constrained optimization problems). When the underlying map $T$ (and by extension the matrix $\mathbf{C}$) is unknown, many natural constraints on the map become linear constraints in its functional representation. In practice, the simplest method for recovering an unknown functional map is to solve the following optimization problem:
@@
where $\mathbf{A}$ and $\mathbf{B}$ are sets of corresponding functions, denoted as descriptors, expressed in the bases on $\mathcal{M}$ and $\mathcal{N}$, respectively, and $\rho(\mathbf{C})$ represents additional constraints deriving from the properties of the matrix $\mathbf{C}$ [45]. When the manifolds $\mathcal{M}$ and $\mathcal{N}$ are approximately isometric and the descriptors are well-preserved by the (unknown) map, this procedure provides a good approximation of the underlying map. In cases where the correspondence $T$ is encoded as a permutation matrix $\mathbf{P}$, the functional map can be retrieved as $\mathbf{C}=\mathbf{\Phi}_{\mathcal{N}}^{{\dagger}}\mathbf{P}\mathbf{\Phi}_{\mathcal{M}}$ where $\mathbf{\Phi}_{\mathcal{M}}$ and $\mathbf{\Phi}_{\mathcal{N}}$ are the bases of the functional spaces $\mathcal{F}(\mathcal{M},\mathbb{R})$ and $\mathcal{F}(\mathcal{N},\mathbb{R})$, respectively, and denotes the Moore-Penrose pseudo-inverse.
@@ 3.2 Latent Functional Maps
3.2.1 Setting
We consider deep neural networks $f:=f_{1}\circ f_{2}\circ...f_{n}$ where each layer $f_{i}$ is associated to a representational space $\mathcal{X}_{i}$ corresponding to the image of $f_{i}$. In the following we’re gonna drop the subscript $i$, when the layer considered is clear form the context. We assume that elements $x\in\mathcal{X}$ are sampled from a latent manifold $\mathcal{M}$. Considering pairs of spaces $(\mathcal{X},\mathcal{Y})$, and corresponding manifolds $\mathcal{M},\mathcal{N}$ our objective is to characterize the relation between them by mapping the space of functions $\mathcal{F}(\mathcal{M})$ to $\mathcal{F}(\mathcal{N})$. Our framework can be summarized in three steps, which we are going to describe in the following sections: (i) how to represent $\mathcal{M}$ from a sample estimate $X$, by building a graph in the latent space $\mathcal{X}$ (ii) how to encode any known preserved quantity between the two spaces by defining a set of descriptor function for each domain (iii) optimizing for the latent functional map between $\mathcal{M}$ and $\mathcal{N}$. An illustrative description of our framework is depicted in Figure 1.
Building the graph.
To leverage the geometry of the underlying manifold, we model the latent space of a neural network building a symmetric k-nearest neighbor (k-NN) graph [31]. Given a set of samples $X=\{{x}_{1},\ldots,{x}_{n}\}$, we construct an undirected weighted graph $G=(X,E,\mathbf{W})$ with nodes $X$, edges $E$, and weight matrix $\mathbf{W}$. The weight matrix is totally characterized by the choice of a distance function $d:\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$.Suitable choices for $d$ include the $L2$ metric or the angular distance. Edges $E$ are defined as $E=\{(x_{i},x_{j})\in X\times X\mid x_{i}\sim_{\text{k}}x_{j}\text{ or }x_{j}\sim_{\text{k}}x_{i}\}$, where $x_{i}\sim_{\text{k}}x_{j}$ indicates that $x_{j}$ is among the $k$ nearest neighbors of $x_{i}$. The weight matrix $\mathbf{W}\in\mathbb{R}^{n\times n}_{\geq 0}$ assigns a weight $\omega(x_{i},x_{j})=\alpha(d(x_{i},x_{j}))$ to each edge $(x_{i},x_{j})\in E$, and $\mathbf{W}_{i,j}=0$ otherwise, with $\alpha$ being some monotonic function. Next, we define the associated weighted graph Laplacian operator $\mathcal{L}_{G}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{W}\mathbf{D}^{-1/2}$, where $\mathbf{D}$ is the diagonal degree matrix with entries $\mathbf{D}_{i,i}=\sum_{j=1}^{n}\mathbf{W}_{i,j}$. $\mathcal{L}_{G}$ is a positive semi-definite, self-adjoint operator [56]), therefore, it admits an eigendecomposition $\mathcal{L}_{G}=\mathbf{\Phi}_{G}\mathbf{\Lambda}_{G}\mathbf{\Phi}_{G}^{T}$, where $\mathbf{\Lambda}_{G}$ is a diagonal matrix containing the eigenvalues, and $\mathbf{\Phi}_{G}$ is a matrix whose columns are the corresponding eigenvectors. The eigenvectors form an orthonormal basis for the space of functions defined on the graph nodes (i.e., $\mathbf{\Phi}_{G}^{T}\mathbf{\Phi}_{G}=\mathbf{I}$). We give comprehensive details on the choice of $k$ in Appendix A.1.
Choice of descriptors.
To define the alignment problem between pair of spaces $\mathcal{X},\mathcal{Y}$ we will introduce the notion of descriptor function. We define as descriptor function any real valued function $f:G\mapsto\mathbb{R}^{k}$ defined on the nodes of the graph. Informally descriptors should encode the information which is shared between $\mathcal{X}$ and $\mathcal{Y}$, either explicitly or implicitly . We categorize descriptor functions into supervised, weakly supervised and unsupervised descriptors. For the former we assumed to have access to a partial correspondence between the domains $\mathcal{X},\mathcal{Y}$, defined as a bijective mapping $\Gamma_{S}:\mathcal{A}_{\mathcal{X}}\mapsto\mathcal{A}_{\mathcal{Y}}$ where $\mathcal{A}_{\mathcal{X}}\subset\mathcal{X},\mathcal{A}_{\mathcal{Y}}\subset\mathcal{Y}$. Then descriptors takes the form of distance functions $d_{\mathcal{X}}:\mathcal{X}\times\mathcal{A}_{\mathcal{X}}\mapsto\mathbb{R}^{+}$, $d_{\mathcal{Y}}:\mathcal{Y}\times\Gamma_{S}(\mathcal{A}_{\mathcal{X}})\mapsto\mathbb{R}^{+}$. As an example considering the geodesic distance (shortest path distance on the k-NN graph) from each node of the graph to the nodes in the anchor set, is an instance of supervised descriptor. In the case of weakly supervised descriptor, we assume to have access to an equivalence relation defined $\Gamma_{W}:\mathcal{A}_{\mathcal{X}}\times\mathcal{A}_{\mathcal{Y}}\mapsto\left\{0,1\right\}$. Example of these are multi-to-multi mappings, like mappings between labels. Unsupervised descriptors are quantities that depends on the topology and the geometry of the graph itself and they don’t rely on any pre-given correspondence or equivalence relation. An example of this is the Heat Kernel Signature [52] node descriptor, based on heat diffusion over the manifold. We give examples of descriptors and ablation in Appendix C.2.
@@ 3.2.1 Setting
Computing LFMs. We now describe the optimization problem to compute a latent functional map between $\mathcal{X}$ and $\mathcal{Y}$. We model each space using a subset of training samples $X=\{x_{1},\ldots,x_{n}\}$ and $Y=\{y_{1},\ldots,y_{n}\}$ and build the k-NN graphs $G_{X}$ and $G_{Y}$ from these samples, respectively. For each graph, we compute the graph Laplacian $\mathcal{L}_{G}$ and derive the first $k$ eigenvalues $\Lambda_{G}$ and eigenvectors $\mathbf{\Phi}_{G}=[\phi_{1},\ldots,\phi_{k}]$, which serve as the basis for the function space defined on the latent spaces.
@@
Given the set of descriptors $\mathbf{F}_{G_{X}}=[f^{G_{X}}_{1},\ldots,f^{G_{X}}_{n_{f}}]$ and $\mathbf{F}_{G_{Y}}=[f^{G_{Y}}_{1},\ldots,f^{G_{Y}}_{n_{f}}]$, we consider the optimization problem defined in Equation 1 and incorporate two regularizers which enforce commutativity for the Laplacian and the descriptors, respectively, with the map $\mathbf{C}$:
$\underset{\mathbf{C}}{\arg\min}\|\mathbf{C}\hat{\mathbf{F}}_{G_{X}}-\hat{\mathbf{F}}_{G_{Y}}\|_{F}^{2}+\alpha\rho_{\mathcal{L}}(\mathbf{C})+\beta\rho_{f}(\mathbf{C})$ (2)
where $\hat{\mathbf{F}}_{G}=\mathbf{\Phi}_{G}^{T}\mathbf{F}_{G}$ are the spectral coefficients of the functions $\mathbf{F}_{G}$, $\rho_{\mathcal{L}}$ and $\rho_{f}$ are the Laplacian and descriptor operator commutativity regularizers respectively, akin to [41]. We give full details on the regularizers in Appendix A. Once we have solved the optimization problem defined in Equation 2, we refine the resulting functional map $\mathbf{C}$ using the spectral upsampling algorithm proposed by [34], as detailed in Appendix A.3.